OpenStudy (anonymous):

Which of the following is a factor of the polynomial x^2-x-20? A. x-5 B. x-4 C. x+2 D. x+5 E. x+10

OpenStudy (anonymous):

do you know how to factor it first?

OpenStudy (anonymous):

I can't for the life of me remember how to do it. I haven't had a math class in roughly two years, and I have to take a placement test for class soon, so I've been doing some prep.

OpenStudy (anonymous):

ok what 2 numbers multiply together =20 and add together =1

OpenStudy (anonymous):

-5 and 4

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&-x&-20\\ &\uparrow&\uparrow \\ &4-5&4\cdot -5 \end{array}\)

OpenStudy (anonymous):

correct. so now you have x^2-x-20=0-----> (x-5) and (x+4) now all u have to do is to check which one will give you 0

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

u got it :D

OpenStudy (anonymous):

you know why "2 numbers multiply together =20 and add together =-1"?